OpenStudy (anonymous):

i will medal and fan Solve 2x + 5 = 27 A.64 B.8.5 C.11 D.16

OpenStudy (anonymous):

So how to solve for a variable (x), you have to get it by itself on one side of the = sign. So you'd first move the 5 over by subtracting it on both sides. Then divide both sides by 2. Like: 2x+5=27 2x=22 x=? Hope this helps!

OpenStudy (anonymous):

C

OpenStudy (shaik0124):

2x+5-5=27-5 2x=22 2x/2=22/2 x=11

OpenStudy (anonymous):

Thanks all of you!

OpenStudy (anonymous):

NP :) tag me if you need any more help

OpenStudy (anonymous):

Okay! :)

OpenStudy (shaik0124):

u said something@kylie :)